# Femtoscopy of $DN$ and $\bar{D}N$ systems

### v1

@@ III.2 TROY framework
In this subsection, we present the results obtained using the TROY formalism including all coupled channels in the off-shell $T$-matrix approach. We also use the ZR approximation with $g=6.0$ but adjust the value of the regulator to $\Lambda=925$ MeV so that the $I=0$ amplitude of the $DN$ channel has a maximum at the position of the $\Lambda_{c}(2595)$ state. Notice that the regulator $\Lambda$ plays an equivalent role to $k_{\textrm{max}}$ in the on-shell calculation, but they do not need to be equal since the off-shell calculation applies it via the form factor of Eq. (9), while the on-shell calculation uses a hard cutoff regularization.
@@
Then, we apply the same regulator for the remaining ${\bar{D}}N$ and $DN$ channels. In particular, in the $I=1$ sector we find a maximum of the $DN$ scattering amplitude at $\sqrt{s}=2798$ MeV, which we identify with the $\Sigma_{c}(2800)$ [62], consistent with the findings of [63].
@@
In the following plots, we compare the TROY results with those of the LL approximation, obtained with the low-energy parameters from the on-shell approach. Again, we use a small Gaussian source radius, $r_{0}=1$ fm, to enhance the effect of the strong interaction in the correlation function.
Figure 3: $\bar{D}^{0}p$ correlation function as function of the relative momentum. We use TROY with the solution of the off-shell $T$-matrix using the ZR approximation potential for the kernel. We use a Gaussian source radius of $r_{0}=1$ fm.
[…]
Figure 4: $D^{-}p$ correlation function as function of the relative momentum. We use TROY with the solution of the off-shell $T$-matrix using the ZR approximation potential for the kernel. We use a Gaussian source radius of $r_{0}=1$ fm.
[…]
The comparison with the LL approximation (short-dashed red line) shows some quantitative differences. While the TROY result remains always below the Coulomb only correlation function, and stays below one, the LL result remains always above one, and crosses the Coulomb only correlation function around $k=120$ MeV.
Figure 5: $D^{0}p$ correlation function as function of the relative momentum. We use TROY with the solution of the off-shell $T$-matrix using the ZR approximation potential for the kernel. We use a Gaussian source radius of $r_{0}=1$ fm.
[…]
Figure 6: $D^{+}p$ correlation function as function of the relative momentum. We use TROY with the solution of the off-shell $T$-matrix using the ZR approximation potential for the kernel. We use a Gaussian source radius of $r_{0}=1$ fm.
@@
Channel $a_{0}$ (fm) ZR TROY $a_{0}$ (fm) ZR on shell
[…]
@@ III.3 Coupled channels in the $D^{+}p$ and $D^{0}p$ correlation functions
In the right panel, we plot the $D^{0}p$ case and its coupled channels, one by one, from the bottom to the top. Here we observe a large effect of the $\pi\Lambda_{c}$ channel, for the same reason as before, that is, because of its coupling to the $I=1$ $\Sigma_{c}(2800)$ resonance and the larger value of the thermal weight. For the $\pi\Sigma_{c}$ case we observe a moderate effect from the $\pi^{+}\Sigma_{c}^{0}$ (difference between the solid green and dashed blue lines) and $\pi^{-}\Sigma_{c}^{++}$ (difference between the solid orange and solid green lines) channels, which are coupled to the $D^{0}p$ state in both the $I=0$ and $I=1$ channels, dominated respectively, by the $\Lambda_{c}(2595)$ and the $\Sigma_{c}(2800)$ resonances. Interestingly, the effect of the $\pi^{0}\Sigma_{c}^{+}$ (difference between the dashed red and blue lines) is almost zero. We note that this state only couples to the $D^{0}p$ via the $I=0$ channel (the $I=1$ has zero Clebsch-Gordan coefficient). We therefore conclude that the relevance of a channel in the $D^{0}p$ correlation function is driven by its coupling to the $\Sigma_{c}(2800)$ resonance, which lies close to the kinematical region of interest. Finally, we observe that the largest coupled channel contribution is due to the $D^{+}n$ channel, which has a threshold very close to the $D^{0}p$ one, and therefore a similar thermal weight. This contribution is enhanced by the fact that the $DN$ coupling to the $\Sigma_{c}(2800)$ resonance is about an order of magnitude larger than that of the $\pi\Lambda_{c}$ and $\pi\Sigma_{c}$ states [63]. The channels with thresholds above the $D^{+}n$ one provide a negligible contribution to the $D^{0}p$ correlation function.
Figure 7: Correlation function of $DN$ for a Gaussian source with $r_{0}=1$ fm. Left panel: $D^{+}p$ with only strong interaction. Right panel: $D^{0}p$. The lines are to be read from bottom to the top when we add additional channels up to complete all of them (“all channels”)
@@ III.4 Dependence on the source size and comparison with data
In Fig. 8 we plot the $D^{-}p$ correlation function using TROY for the meson-baryon interaction in the ZR approximation. It represents an overall (Coulomb plus strong) attractive interaction, which cannot fully explain the current experimental data by ALICE [27]. We have used a Gaussian radius between 0.8 fm and 1.0 fm, relevant for pp collisions, and present the results with a blue band. In Ref. [27] only the parametrizations by Haidenbauer—with an attractive interaction in the strong sector given by an additional strength of the scalar exchange—and Yamaguchi, with a bound state, show a reasonable agreement with the data. In the Yamaguchi case, as can be seen in our Fig. 2, we obtain a result consistent with the presence of the bound state. We have not been able to reproduce the curve shown in the ALICE paper [27] for this model, unless one naively assumes $a_{D^{-}p}\simeq a_{\bar{D}N,I=0}$.
Figure 8: $D^{-}p$ correlation function using TROY solution of the ZR approximation. The Gaussian radii is varied between $r_{0}=(0.9\pm 0.1)$ fm, to cover small systems as those measured by ALICE in high-multiplicity pp collisions.
@@
In Fig. 9 we present the results of the $D^{0}p$ and $\bar{D}^{0}p$ channels using the full coupled-channel TROY formalism with thermal weights. The Gaussian radii are chosen to be $r_{0}=1$ fm, 3 fm, and 5 fm. The latter is used to mimic the source sizes expected in Au+Au collisions measured by the STAR collaboration. The signal for these large radii is rather weak and suppressed, but our result is roughly consistent with the preliminary data shown in Ref. [64]. However, in that reference the $D^{0}p$ and $\bar{D}^{0}p$ cases are summed up to a single correlation function. Since the interaction is rather different for each of these sectors, we doubt this is a sensible comparison and urge our experimental colleagues to provide each channel separately.
@@ IV Summary
In this work we have studied the femtoscopic correlation function of $D^{+}p,D^{0}p,D^{-}p$ and $\bar{D}^{0}p$ pairs applying the Koonin-Pratt equation. The meson-baryon interaction has been extracted from an effective field theory at low energy, working in the zero-range approximation. Non perturbative $T$-matrix equations have been implemented both in the on-shell factorization approach, as well as in the off-shell case using the TROY formalism. These two methods provide different approximations for the $l=0$ wave function of the meson-baryon pair, which eventually are used in the Koonin-Pratt formula. Coulomb effects have been also incorporated in the appropriate channels.
@@
The on-shell calculation was used to extract low-energy parameters of the interaction which were employed in the Lednický-Lyuboshitz approximation of the correlation function. We remind the reader that this approximation is based on the asymptotic behavior of the wave function of the direct channel, thereby neglecting the effect of the inelastic components. For the $D^{+}p$ case we found that different results in the literature (with the exception of the SU(8) calculation) provide a similar correlation function reflecting a repulsive strong interaction. For the $D^{-}p$ pair we also find a consistency among the different results despite the different values of the scattering lengths, with a clear difference of the Haidenbauer’s calculation including an enhanced $\sigma$ exchange which encodes a rather strong attraction. This result was close to the experimental points as shown in Ref. [8]. We stress that in our version of the LL formula, the Yamaguchi result does not approach the data, but shows the typical effect of a bound state, in accordance with the theoretical expectations.
@@
The full TROY calculation incorporates all the inelastic channels and obtains the complete spatial dependence of the wave function (not just the asymptotic part) out of the off-shell $T$-matrix elements. Moreover, we have estimated the thermal weights for all channels taking into account the masses of the produced pairs in a system at finite temperature for conditions given at high-multiplicity pp collisions at $\sqrt{s}=13$ TeV in the LHC. For the $D^{+}p$ case we found in our model that the direct channel is roughly compatible with the LL approximation, but the effect of the coupled channels is non negligible, increasing the correlation function between 30-50% at low momentum. We provide this correlation function for a small source radius as a prediction for future measurements of this system in the ALICE experiment, in which we foresee that the LL approach will not be enough. For the $D^{-}p$ case, the effect of the only coupled channel $(\bar{D}^{0}n)$ is found to minimal with respect to the direct one, but there are differences between the LL and TROY calculations of the correlation function, the latter maintaining values above one for all momenta, while the former approaches one from below. We find no consistency with the experimental data for $D^{+}p$ pairs measured by ALICE in pp collisions. Finally, we have computed the correlation function for neutral channels, $D^{0}p$ and $\bar{D}^{0}p$, motivated by the preliminary results by the STAR collaboration. Since these correspond to Au+Au collisions we scanned large Gaussian radii between $r_{0}=1$ fm and $r_{0}=5$ fm. We have not quantitatively compared with preliminary data since they are given as the sum of the two channels, and we cannot really assess the different nature of the interaction in the $C=1$ and $C=-1$ sectors.
@@
To further test the results of this paper, it would be vital to compare the $D^{+}p$ correlation function in small systems like pp collisions, for which we present a prediction but no experimental results have been provided yet by the ALICE experiment. This channel could be an excellent way to study the effect of the $\Sigma_{c}(2800)$ resonance. In addition, it would be interesting to compare with STAR results in larger systems, like the one shown in [64], but after the data of the two charm sectors are given separately. While this would lead to an increase of the statistical uncertainties, it would make the comparison more sensible, since in our model the $D^{0}p$ and $\bar{D}^{0}p$ interactions are rather different, namely, strongly attractive and repulsive, respectively.

### v2

@@ III.2 TROY framework
In this subsection, we present the results obtained using the TROY formalism including all coupled channels in the off-shell $T$-matrix approach. We employ the ZR kernel interaction with $g=6.0$ but adjust the cut-off value to $k_{\textrm{max}}=630$ MeV so that the $I=0$ amplitude of the $DN$ channel has a maximum at the position of the $\Lambda_{c}(2595)$ state. Then, we apply the same regulator for the remaining $DN$ and ${\bar{D}}N$ channels. In particular, in the $I=1$ sector we find a maximum of the $DN$ scattering amplitude at $\sqrt{s}=2798$ MeV, which we identify with the $\Sigma_{c}(2800)$ [63], consistent with the findings of Ref. [64].
@@
In the following plots, we compare the TROY results (solid black lines) with those of the LL approximation (short-dashed red lines), obtained with the corresponding TROY low-energy parameters (first column in Table 3). We also show results for the correlation function obtained with the on-shell version of the $T$-matrix (dash-dotted blue lines). Again, we use a small Gaussian source radius, $r_{0}=1$ fm, to enhance the effect of the strong interaction in the correlation function.
Figure 3: $\bar{D}^{0}p$ correlation function as function of the relative momentum for a source size of $r_{0}=1$ fm. We show the full TROY result with the ZR kernel (solid black line), the LL model (short-dashed red line), and the correlation function obtained with the on-shell version of the $T$-matrix (dash-dotted blue line).
We start with the simple $\bar{D}^{0}p$ correlation function in Fig. 3. This is the unique channel in the C=1,S=0,Q=1C=-1,S=0,Q=1 sector. The TROY correlation function is very similar to the on-shell case and to the LL result, showing the effect of a repulsive interaction as expected from the negative value of the scattering length, $a_{0}=(-0.35+\rm{i}0)$ fm.
Figure 4: $D^{-}p$ correlation function as function of the relative momentum for a source size of $r_{0}=1$ fm. We show the full TROY result with the ZR kernel (solid black line), the elastic contribution to the TROY result (dashed green line), the LL model (short-dashed red line), and the correlation function obtained with the on-shell version of the $T$-matrix (dash-dotted blue line). The dotted blue line displays the Coulomb-only result.
In Fig. 4 we present the $D^{-}p$ correlation function. The full TROY result contains the effect of the $D^{-}p$ and $\bar{D}^{0}n$ coupled channels present in this C=1,S=0,Q=0C=-1,S=0,Q=0 sector. The inelastic process $\bar{D}^{0}n\rightarrow D^{-}p$ contributes with a thermal weight very close to one and its effect is extremely small, as can be seen by the difference between the solid black line and the elastic contribution represented by the long-dashed green line. The elastic channel has a scattering length of $a_{0}=(-0.16+\rm{i}0.01)$ fm, signalling a repulsive strong interaction on top of the attractive Coulomb one, the correlation function of which is represented by the dotted blue line. The comparison with the LL approximation shows some quantitative differences. While the TROY result remains always below the Coulomb only correlation function, and stays below one, the LL result remains always above one, and crosses the Coulomb only correlation function around $k=120$ MeV.
Figure 5: $D^{0}p$ correlation function as function of the relative momentum for a source size of $r_{0}=1$ fm. We show the full TROY result with the ZR kernel (solid black line), the elastic contribution to the TROY result (dashed green line), the LL model (short-dashed red line), and the correlation function obtained with the on-shell version of the $T$-matrix (dash-dotted blue line).
Moving to the charm one sector, we start by showing in Fig. 5 the correlation function of $D^{0}p$ pairs, free from the Coulomb interaction and involving 16 coupled channels. The full TROY result contains the contribution of all these channels with some thermal weights, giving a large contribution at low momentum values (4 channels are below the elastic one, and 11 channels above it, see Sec. III.3). The cusp observed at $k=88$ MeV signals the opening of the $D^{+}n$ channel. In this case the TROY and on-shell results differ more significantly, as the interaction is now sensitive to the existence of a resonance below but very close to threshold, the $\Sigma_{c}(2800)$ in $I=1$. This is reflected in the large negative real part of the $D^{0}p$ scattering length, which is $a_{0}=(-1.62+\rm{i}0.48)$ fm in the TROY formalism, to be compared with the on-shell value of $a_{0}=(-1.44+\rm{i}0.35)$ fm. Note that the $D^{0}p$ interaction also develops a bound state with $I=0$, the $\Lambda_{c}(2595)$. This state should leave a visible effect in the correlation functions of $\Sigma_{c}\pi$ pairs, which are not the object of the present study. The elastic contribution, represented by the long-dashed green line, also differs considerably from the LL result, displayed by the short-dashed red line. This has to do with the short-distance behavior of the wave function, which differs significantly from the asymptotic form assumed by the LL approach. This is especially significant at low momentum, since the wave function is more affected by the influence of the subthreshold resonance.
Finally, we focus on the $D^{+}p$ correlation as shown in Fig. 6. This sector is affected by the Coulomb interaction and has 9 coupled channels. Again, the direct contribution to the correlation function is quite comparable to the LL result, except at low momentum as explained above. Inelastic channels provide a sizable contribution, especially around $k=50$ MeV, as will be detailed in Sec. III.3, probing the importance of coupled channels in the $D^{+}p$ case, where the LL approximation would not be appropriate. The difference between the TROY and on-shell results can be understood by their different scattering lengths in the elastic $D^{+}p$ channel, namely $a_{0}=(-1.97+\rm{i}0.39)$ fm and $a_{0}=(-1.79+\rm{i}0.32)$ fm, respectively. As before, these sizable values are directly related to the $I=1$ $\Sigma_{c}(2800)$ resonance present in both models.
Figure 6: $D^{+}p$ correlation function as function of the relative momentum for a source size of $r_{0}=1$ fm. We show the full TROY result with the ZR kernel (solid black line), the elastic contribution to the TROY result (dashed green line), the LL model (short-dashed red line), and the correlation function obtained with the on-shell version of the $T$-matrix (dash-dotted blue line). The dotted blue line displays the Coulomb-only result.
@@
Channel $a_{0}$ (fm) ZR TROY $a_{0}$ (fm) ZR on shell (charge) $a_{0}$ (fm) ZR on shell (isospin)
$\bar{D}^{0}p$ $0.35+\rm{i}0$ $0.34+\rm{i}0$ $0.34+\rm{i}0$
$D^{-}p$ $0.16+\rm{i}0.01$ $0.16+\rm{i}0.01$ $0.16+\rm{i}0$
$D^{0}p$ $1.62+\rm{i}0.48$ $1.44+\rm{i}0.35$ $1.21+\rm{i}0.16$
$D^{+}p$ $1.97+\rm{i}0.39$ $1.79+\rm{i}0.32$ $1.79+\rm{i}0.32$
@@
Before concluding this section, we would like to comment on the concerns raised in Ref. [65] about the use of a universal source function and the resulting ambiguities in the correlation function tied to the off-shell behavior of the wave-function or, equivalently, to the off-shell dependence of the $T$ matrix. We indeed find that the way the off-shell treatment is done, such as the type of form-factor employed, or the value of the cut-off, influences the results, especially in the $DN$ sector. However, using physically motivated parameters, fine tuned to experimental constraints (like the position of the generated resonances) help to minimize this dependence. In fact, the interactions used in the present work can be thought of being built by the exchange of vector mesons and a cut-off of the order of the vector mass (6001000600-1000 MeV), as selected here, should be appropriate. Besides, the study on Ref. [66] explicitly shows, for a meson-baryon interaction of the type of the present work but in the strangeness $S=0$ sector around the $N^{*}(1535)$ resonance, that the uncertainties in the correlation functions due to off-shell ambiguities are of the order of 2–3%.
@@ III.3 Coupled channels in the $D^{+}p$ and $D^{0}p$ correlation functions
In the right panel, we plot the $D^{0}p$ case and its coupled channels, one by one, from the bottom to the top. Here we observe a large effect of the $\pi\Lambda_{c}$ channel, for the same reason as before, that is, because of its sizable coupling to the $I=1$ $\Sigma_{c}(2800)$ resonance and the larger value of the thermal weight. For the $\pi\Sigma_{c}$ case we observe a moderate effect from the $\pi^{+}\Sigma_{c}^{0}$ (difference between the solid green and dashed blue lines) and $\pi^{-}\Sigma_{c}^{++}$ (difference between the solid orange and solid green lines) channels, which are coupled to the $D^{0}p$ state in both the $I=0$ and $I=1$ channels, dominated respectively, by the $\Lambda_{c}(2595)$ and the $\Sigma_{c}(2800)$ resonances. Interestingly, the effect of the $\pi^{0}\Sigma_{c}^{+}$ (difference between the dashed red and blue lines) is almost zero. We note that this state only couples to the $D^{0}p$ via the $I=0$ channel (the $I=1$ has zero Clebsch-Gordan coefficient). We therefore conclude that the relevance of a channel in the $D^{0}p$ correlation function is driven by its coupling to the $\Sigma_{c}(2800)$ resonance, which lies close to the kinematical region of interest. Finally, we observe that the largest coupled channel contribution is due to the $D^{+}n$ channel, which has a threshold very close to the $D^{0}p$ one, and therefore a similar thermal weight. This contribution is enhanced by the fact that the $DN$ coupling to the $\Sigma_{c}(2800)$ resonance is about an order of magnitude larger than that of the $\pi\Lambda_{c}$ and $\pi\Sigma_{c}$ states [64]. The channels with thresholds above the $D^{+}n$ one provide a negligible contribution to the $D^{0}p$ correlation function.
Figure 7: Correlation function of DNDN for a Gaussian source with $r_{0}=1$ fm. Left panel: $D^{+}p$ with only strong interaction. Right panel: $D^{0}p$. The lines are to be read from bottom to the top when we add additional channels up to complete all of them (“all channels”)
@@ III.4 Dependence on the source size and comparison with data
In Fig. 8 we plot the $D^{-}p$ correlation function using TROY for the meson-baryon interaction in the ZR approximation. It represents an overall (Coulomb plus strong) attractive interaction, which cannot fully explain the current experimental data by ALICE [28]. We note that the blue band displaying our results collects the correlation functions of 2000 configurations obtained by varying the Gaussian source radius between 0.8 fm and 1.2 fm, relevant for pp collisions, the coupling constant $g$ between 4.5 and 7.5, and the hard cut-off $\Lambda$ between 550 MeV and 750 MeV. Even with that flexibility in the model parameters we are unable to reproduce the current ALICE data. In Ref. [28] only the parametrizations by Haidenbauer—with an attractive interaction in the strong sector given by an additional strength of the scalar exchange—and Yamaguchi, with a bound state, show a reasonable agreement with the data. In the Yamaguchi case, as can be seen in our Fig. 2, we obtain a result consistent with the presence of the bound state.
Figure 8: $D^{-}p$ correlation function using TROY solution of the ZR approximation. The Gaussian source radius is varied between 0.8 fm and 1.2 fm, relevant for pp collisions, the coupling constant $g$ between 4.5 and 7.5, and the hard cut-off $\Lambda$ between 550 MeV and 750 MeV radii is varied between $r_{0}=(0.9\pm 0.1)$ fm.
@@
In Fig. 9 we present the results of the $D^{0}p$ and $\bar{D}^{0}p$ channels using the full coupled-channel TROY formalism with thermal weights. The Gaussian radii are chosen as $r_{0}=1$ fm, 3 fm, and 5 fm. The latter is used to mimic the source sizes expected in Au+Au collisions measured by the STAR collaboration. The signal for these large radii is rather weak and suppressed, but our result is roughly consistent with the preliminary data shown in Ref. [67] for the larger radii. However, in that reference, the $D^{0}p$ and $\bar{D}^{0}p$ cases are combined into a single correlation function, even if they are not charge conjugate channels. Since the interaction is rather different for each of these sectors, we doubt that this is a sensible comparison and urge our experimental colleagues to provide each channel separately.
@@ IV Summary
In this work we have studied the femtoscopic correlation function of $D^{+}p,D^{0}p,D^{-}p$ and $\bar{D}^{0}p$ pairs applying the Koonin-Pratt equation. The meson-baryon interaction has been extracted from an effective field theory at low energy, working in the zero-range approximation. Non perturbative $T$-matrix equations have been implemented both in the on-shell factorization approach, as well as in the off-shell case using the TROY formalism. These two methods provide different approximations for the $l=0$ wave function of the meson-baryon pair, which eventually are used in the Koonin-Pratt formula. Coulomb effects have also been incorporated in the appropriate channels. Our results have been compared to those obtained within the Lednický-Lyuboshitz approximation of the correlation function. We remind the reader that this approximation is based on the asymptotic behavior of the wave function of the direct channel, thereby neglecting the effect of the inelastic components as well as possible deviations of the wave function at small distances, which can be important at low momentum if a nearby subthreshold resonance is present. For the $D^{+}p$ case we found that,(with the exception of the SU(8) calculation, different results in the literature provide a similar correlation function reflecting an apparently repulsive strong interaction. For the $D^{-}p$ pair we also find consistency among the different results despite the different values of the scattering lengths, with a clear difference of the Haidenbauer’s calculation including an enhanced $\sigma$ exchange which encodes a rather strong attraction. This result was close to the experimental data as shown in Ref. [9]. We stress that in our version of the LL formula, the Yamaguchi result does not approach the data, but shows the typical effect of a bound state, in accordance with the theoretical expectations.
@@
The full TROY calculation incorporates all the inelastic channels and obtains the complete spatial dependence of the wave function (not just the asymptotic part) out of the off-shell $T$-matrix elements. Moreover, we have estimated the thermal weights for all channels, taking into account the masses of the produced pairs in a system at finite temperature for the conditions of high-multiplicity pp collisions at $\sqrt{s}=13$ TeV in the LHC. For the $D^{+}p$ case we found the direct elastic channel of our model to be roughly compatible with the LL approximation, but the effect of the coupled channels is non negligible, increasing the correlation function between 30-50% at low momentum. We provide this correlation function for a small source radius as a prediction for future measurements of this system in the ALICE experiment, in which we foresee that an analysis based on the LL approach will not be appropriate enough. For the $D^{-}p$ case, the effect of the only coupled channel $(\bar{D}^{0}n)$ is found to be minimal with respect to the direct one, but there are differences between the LL and TROY calculations of the correlation function, the latter maintaining values above one for all momenta, while the former approaches one from below. We find no consistency with the experimental data for $D^{+}p$ pairs measured by ALICE in pp collisions. Finally, we have computed the correlation function for neutral channels, $D^{0}p$ and $\bar{D}^{0}p$, motivated by the preliminary results by the STAR collaboration. Since these correspond to Au+Au collisions we scanned large Gaussian radii between $r_{0}=1$ fm and $r_{0}=5$ fm. We have not quantitatively compared with preliminary data since they are given as the sum of the two channels, and we cannot really assess the different nature of the interaction in the $C=1$ and C=1C=-1 sectors.
@@
To further test the results of this paper, it would be vital to compare with the $D^{+}p$ correlation function in small systems like pp collisions, for which we present a prediction, but no experimental results have been yet provided by the ALICE experiment. This channel could be an excellent way to study the effect of the $\Sigma_{c}(2800)$ resonance. In addition, it would be interesting to compare with STAR results in larger systems, like the one shown in [67], but after the data of the two charm sectors are given separately. While this would lead to an increase of the statistical uncertainties, it would make the comparison more sensible, since in our model the $D^{0}p$ and $\bar{D}^{0}p$ interactions are rather different, namely, strongly attractive and repulsive, respectively.